Need to find the equations for x in this word problem. Thanks!
A projectile is launched vertically upward with an initial speed of 256 ft/sec. The equation for its vertical distance from the ground is h = 256t – 16t^2, where t is time in seconds. Find the projectile’s maximum height and the total time that it is in the air.

Answer by fractalier(2101)

Remember that an object in upward flight is slowed by gravity at a rate of -32 ft/s/s...
Thus if an object has an initial upward velocity of 256 ft/s, it reaches a speed of zero after 8 seconds. (256 - 8(32) = 0)
Now, knowing this, we can plug t = 8 into the height function and get
h = 256t – 16t^2
h = 256(8) - 16(8^2)
h = 2048 - 1024
h = 1024 feet